### v2

@@ E.4   Revisiting the ”issue” regarding the use of the Wald-DID estimand
Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand in Fuzzy DID designs, and show that while their argument is correct, it is misguided. During this discussion, we also clarify why dCDH exclude so many types in exposed and unexposed groups as we see in Appendix E.2. Hereafter, we assume Assumptions 24 - 28 in Fuzzy DID designs.
dCDH argue that the Wald-DID estimand requires Assumption 28 (stable treatment effect over time) in order to identify the SLATET, and propose alternative estimands that do not rely on this assumption. We first show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome, which is common in DID designs, instead of on the parallel trends assumption in the unexposed outcome as in this article. We next argue that the former type of parallel trends assumption is unsuitable for DID-IV settings.
First, we clarify the interpretation of Assumption 28 by introducing the additional notation. This assumption is actually being conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but dCDH do not provide the reason why this strong restriction is required only for the sub-population. For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$. Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
Then, from Table 9 and Table 10, we can decompose Assumption 28 in the following:
@@
where we use the Law of Iterated Expectation, $D_{0}(0)=D_{0}(1)$ (implied by the sharp assignment of the instrument), and $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Equation (44) shows that in an unexposed group ($z=0$), Assumption 28 requires the stable treatment effect in the time always-takers $AT^{T}$. Equation (45) shows that in an exposed group ($z=1$), this assumption requires that a weighted average of the treatment effects among the time defiers $DF^{T}$ and the time always takers $AT^{T}$ is stable during the two periods.
@@
To see this formally, we use our notation to briefly sketch the proof of dCDH on the Wald-DID estimand identifying SLATET under fuzzy DID designs. First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand. We can decompose this quantity into two terms:
@@
Similarly, one can decompose the conditional expectation $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$ as follows:
@@
where the type SW (switcher) includes the units who switch from non treatment to treatment during the two periods. The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ in an exposed group, and we should add and subtract this quantity to construct the second term. The first term disappears if we assume condition (45). Here, we note that we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term.
Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 25 and Assumption 26, the Wald-DID estimand captures the SLATET under Assumption 27 (parallel trends assumption in the untreated outcome) and conditions (44) and (45). From the discussions so far, one can conclude that the parallel trends assumption in the untreated outcome leads dCDH to impose conditions (44) and (45).
Here, we add few remarks to clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units in order for the Wald-DID estimand to identify the SLATET. It turns out that this also stems from their reliance on the parallel trends assumption in the untreated outcome.
@@
dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 9-10). This identification result seems not intuitive because the identification variation comes from the policy shock (instrument) in period $1$.
@@
So far, we have confirmed that dCDH’s argument rests on the parallel trends assumption in the untreated outcome. Finally, we argue that this assumption is unsuitable for DID-IV settings for three reasons.
First, this assumption is not sufficient to capture the average time trends of the outcome even in an unexposed group; in DID-IV settings, the units are allowed to take the treatment in the absence of the instrument, and thus we can not observe the untreated outcomes for these types in an unexposed group.
Second, in most of the DID-IV applications, we can not impute the untreated potential outcomes in general. For instance, Black et al. (2005) estimates the causal link between parents’ and children’s education attainment, exploiting variation in the timing of the implementation of the school reform across municipalities and cohorts as an instrument for parents’ education attainment. In the author’s setting, the parallel trends assumption in the untreated outcome requires that the data contain the parents with zero education attainment, which is unrealistic in practice.
Finally, this assumption is generally not testable using the pre-exposed period data in a given application. This seems apparent because some units can already adopt the treatment before period 00, and we can not compare the pretrends of the untreated outcome between exposed and unexposed groups.
Although dCDH propose the placebo test in their appendix, we show that it is incomplete for checking the plausibility of the parallel trends assumption in the untreated outcome. dCDH’s test consists of two steps. First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$:
@@
Although dCDH ”assume” the conditions (46) and (47), dCDH also note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 27. We therefore describe the dCDH’s test in two steps.
dCDH’s test has some drawbacks. First, the conditions (46) and (47), which we should check in the first step, seem strong, but dCDH do not explain how we should proceed when these conditions are violated in practice. Second, in this test, the second step does not directly test Assumption 27 in the pre-exposed periods; it does not compare the evolution of the mean untreated outcome between exposed and unexposed group before period $T=0$. Indeed, dCDH describe the equality (48) as a necessary condition for Assumption 27 and Assumption 28 to be satisfied under the conditions (46) and (47). Finally, it is unclear how we should conduct this test when we have multiple pre-exposed period data. Although dCDH assume ”for instance that data is available for the period $T=-1$”, they do not provide any guidance on how to perform this test in multiple pre-exposed period settings.
Surprisingly, dCDH’s test corresponds to our pretrends test in section 5.3 in this article if we replace the conditions (46)- (47) in the first step with the weaker one:
@@
Under repeated cross section settings, one can figure out that the conditions (48) - (49) coincide with the conditions (3) -(4) in section 5.3 in this article.

### v4

@@ E.4   Revisiting the ”issue” regarding the use of the Wald-DID estimand
Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand in Fuzzy DID designs, and show that their argument is misguided. We also clarify why dCDH exclude so many types in exposed and unexposed groups as we see in Appendix E.2. Hereafter, we assume Assumptions 24 - 28 in Fuzzy DID designs.
dCDH argue that the Wald-DID estimand requires Assumption 28 (stable treatment effect over time) to identify the SLATET, and propose alternative estimands that do not rely on this assumption. We show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome, which is common in DID designs, instead of on the parallel trends assumption in the unexposed outcome as in this article. Next, we argue that the former type of parallel trends assumption is unsuitable for DID-IV settings.
First, we clarify the interpretation of Assumption 28 by introducing the additional notation. This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but dCDH do not provide the reason why this strong restriction is required only for the sub-population. For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$. Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
Then, from Table 9 and Table 10, we can decompose Assumption 28 as follows:
@@
where we use the Law of Iterated Expectation, $D_{0}(0)=D_{0}(1)$ (implied by the sharp assignment of the instrument), and $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Equation (44) shows that in an unexposed group ($z=0$), Assumption 28 requires the stable treatment effect for the time always-takers $AT^{T}$. Equation (45) shows that in an exposed group ($z=1$), this assumption requires that a weighted average of the treatment effects among the time defiers $DF^{T}$ and the time always takers $AT^{T}$ is stable during the two periods.
@@
To see this formally, we use our notation to briefly sketch the proof of dCDH on the Wald-DID estimand identifying SLATET under fuzzy DID designs. First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand. We can decompose this quantity into two terms.
@@
Similarly, one can decompose the conditional expectation $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$ as follows.
@@
where the type SW (switcher) includes the units who switch from non-treatment to treatment during the two periods. The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ in an exposed group, and we should add and subtract this quantity to construct the second term. The first term disappears if we assume condition (45). Here, we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term.
Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 25 and Assumption 26, the Wald-DID estimand captures the SLATET under Assumption 27 (parallel trends assumption in the untreated outcome) and conditions (44) and (45). From the discussions, one can conclude that the parallel trends assumption in the untreated outcome leads dCDH to impose conditions (44) and (45).
Here, we clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units for the Wald-DID estimand to identify the SLATET. This also stems from their reliance on the parallel trends assumption in the untreated outcome.
@@
dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 9-10). This identification result does not seem intuitive because the identification variation comes from the policy shock (instrument) in period $1$.
@@
We have confirmed that dCDH’s argument regarding the use of the Wald-DID estimand rests on the parallel trends assumption in the untreated outcome in DID designs. Finally, we argue that this assumption is unsuitable for DID-IV settings for three reasons.
First, this assumption is not sufficient to capture the average time trends of the outcome even in an unexposed group. This is because in DID-IV settings, the units are allowed to take the treatment in the absence of the instrument, and we cannot observe the untreated outcomes for these types in an unexposed group.
Second, in most of the DID-IV applications, we can not impute the untreated potential outcomes in general. For instance, Black et al. (2005) estimates the causal link between parents’ and children’s education attainment, exploiting variation in the timing of the implementation of the school reform across municipalities and cohorts as an instrument for parents’ education attainment. In the authors’ setting, the parallel trends assumption in the untreated outcome requires the data to contain parents with zero education attainment, which is unrealistic and not what the authors would have in mind in practice.
Finally, this assumption is generally not testable using the pre-exposed period data in a given application. This is apparent, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
dCDH propose the placebo test in their appendix, but we show that it is not enough for checking the plausibility of the parallel trends assumption in the untreated outcome. dCDH’s test has two steps. First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$.
@@
Although dCDH ”assume” the conditions (46) and (47), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 27. We, therefore, describe the dCDH’s test in two steps.
dCDH’s test has some drawbacks. First, the conditions (46) and (47), which should be checked in the first step, are strong, but dCDH do not explain the way ahead when these conditions are violated in practice. Second, in this test, the second step does not directly test Assumption 27 in the pre-exposed periods, as it does not compare the evolution of the mean untreated outcome between exposed and unexposed groups before period $T=0$. Indeed, dCDH describe the equality (48) as a necessary condition for Assumption 27 and Assumption 28 to be satisfied under the conditions (46) and (47). Finally, it is unclear how this test should be conducted with multiple pre-exposed period data. dCDH assume ”for instance that data is available for the period $T=-1$”, but they do not indicate how this test can be performed in multiple pre-exposed period settings.
Surprisingly, dCDH’s test corresponds to our pre-trend test in Subsection 5.3 if we replace the conditions (46)- (47) in the first step with the weaker one.
@@
Under repeated cross section settings, the conditions (48) - (49) coincide with the conditions (3) -(4).